

********************************************************************
UARS WEEKLY STATUS REPORT - 08 OCT 2004
********************************************************************
***********
SOLSTICE
***********
This week, SOLSTICE continued to operate in good health gathering solar
data on all available orbital tracking periods. On average, we received
64% of each day's data amounting to about 570 minutes of solar
observing time per day.
***********
SUSIM
***********
During this past week and as available telemetry time permitted,
SUSIM successfully gathered daily mid (1.1 nm) resolution, MgII,
occultation, solar pointing offset, and continuous monitoring
scans.
***********
PEM
***********
HEPSA data files are version 2. Currently, HEPSA data between 1991274 and
2004139 has been reprocessed and archived at the GSFC DAAC. HEPSA Level 2
data files from 2004140 to 2004159 are processed and waiting for archiving.
MEPS data files are version 1. Processed MEPS data files at the DAAC archive
are between 1991274 and 2004139. MEPS Level 2 data files from 2004140 to
2004159 are processed and waiting for archiving.
PEM continues to receive NMC data files from NOAA. No problems have
been encountered.
The PEM data from the GUVI conjunction periods may be found at
http://wwwpem.space.swri.edu/guvi-pem.html. The additional electron data for
the October and November storms of 2003 has been moved and is now located at
http://climatology.space.swri.edu/uars_fromRudy.html. For those users of
SDDAS, update your UARS Meta data and these data times will appear
with full functionality. Proton data from these storms are now located at
http://climatology.space.swri.edu/UARS_images_from_Rudy_(Apr._2004).html
Requests have been received for storm spectral data. Overview plots were
produced and these plots may be found on the web at
http://climatology.space.swri.edu/UARS_images_postedJune2004.html
for the storms of October 29 of 1991, November 8-9 of 1991, April 6-7 of 2000,
March 31 - April 1 of 2001, April 11-12 of 2001, November 6-7 of 2001, and
November 11-12 of 2001.
*Beginning of power orbit day (PEM activation): array current >5 amps, array
voltage >53.5 volts, and battery current >2.0 amps.
End of power orbit day (PEM deactivation): time in power orbit day >4 minutes,
array current <9 amps, and battery currents <0.2 amps.
Power orbit day variation: ~39-40 min at yaw to ~50-51 min at maximum beta.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
MPG
***********
The MPG continues to serve as the point-of-contact for UARS
instrument operations in support of the Science Traceability
Mission of UARS.
Daily processing of instrument Daily Activity Plans continues
with no problems.
***********
CDHF
***********
Nothing received.
***********
FOT
***********
REPORTING PERIOD
This report covers 25 September 2004 (Orbit 71361, GMT Day 269)
through 01 October 2004 (Orbit 71465, GMT Day 275).
SPACECRAFT OPERATION
The observatory is now in Reverse Flight (Northern Hemisphere viewing)
and is performing nominally under two-battery operations. The
instrument operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
SOLSTICE 271/1247 71398 SOLSTICE TO FULLTIME OPS
SUSIM 271/1511 71400 POWER ON
HALOE 273/1324 71429 POWER ON
PEM 274/1321 71444 ZEPS & NEPS TO DAY ONLY
A Forward-to-Reverse Yaw Around maneuver (number 133) was successfully
executed in Orbit 71370 beginning 269/ 15:10:32Z and completing at
269/ 1 5:47:42Z on 25 September 2004. All instruments were in their
off or standby mode (except for SOLSTICE in day-only). Due to the SA
position at 269 degrees, the effective night length of the spacecraft
after the maneuver was approximately 57.62 minutes. Post maneuver
analysis of data indicate that during the Yaw Around, the end-of-night
load bus voltage was 26.08 volts and the average depths of discharge
(DODs) on Batteries 2 and 3 were approximately 20.62 and 19.92 percent
respectively. The next (reverse-to-forward) YAW is scheduled for 02
November 2004 (DOY 307).
The beta angle ranged between the angles of 0.7 degrees to 28.6
degrees for this report period. The beta angle is now increasing
towards a maximum angle of 65 degrees on 13 October 2004 (DOY 287).
Spacecraft battery 1, 2, and 3 performance monitoring continues.
Battery 1 remains OFF the charge relay. Battery 2 and Battery 3 are
currently maintaining greater than 23.6 V end-of-night (EON) load bus
voltage. Battery temperatures are stable with a temperature delta
between Battery 2 and Battery 3 of 0.64 to 0.65 degrees C. Additional
battery performance data for each battery on 24 September and 01
October 2004 are provided in Appendix D.
The clock error ranged from 11.2 to -10.4 msec during this report
period. Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT
for this report. A summary is listed below and the details of ALERT
and Out-Of-Limits occurrences are listed in Appendix B.
OUT-OF-LIMITS = 00
ALERTS = 00
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
FOT = 04
SOLSTICE = 01
SUSIM = 01
HALOE = 02
PEM = 01
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
FOT = 05
Three (3) Anomaly Report remained OPEN at the end of this report
period (see Appendix E).
GROUND SYSTEM OPERATION
Five (5) operational support problems occurred this period. Anomaly
Reports are included in Appendix E.
The data loss calculations are no longer being supplied by Data
Capture Facility (DCF).
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
A problem was found this week in the new release of the CMS code. The
issue is believed to be with the namelist files, but software
development is looking into the exact cause. FDF has also been asked
to provide a User Antenna Views file which contains SMA codes. This
is needed to complete the CMS processing of SMA supports. A meeting
will be held with FDF to discuss this issue. These two problems have
delayed the release of the new CMS code to operations. No new date
has been set for completion.
MEETINGS
-None this report period.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for 02 November 2004 (DOY 307).
This maneuver will be a Reverse to Forward Flight maneuver (Southern
Hemisphere viewing).
Attachments: Appendix A, Real-time Command Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 01 October 2004
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ ------
71370 269/1503 04-275 MACS YAW MANEUVER - BEGIN
71370 269/1615 04-275 MACS YAW MANEUVER - END
71374 269/2204 04-276 OBC/CLK CLK ADJ FR -20.6 TO 0.0 MSEC/DAY
71398 271/1247 04-277 SOLSTICE SOLSTICE TO FULLTIME OPS
71400 271/1511 04-278 SUSIM BEGIN POWER ON
71411 272/0921 04-278 SUSIM POWER ON COMPLETE
71429 273/1324 04-279 HALOE POWER ON - PART I
71431 273/1644 04-280 HALOE POWER ON - PART II
71444 274/1321 04-281 PEM ZEPS & NEPS TO DAY ONLY
71449 274/2118 04-282 OBC/CLK CLK ADJ FR 0.0 TO -20.6 MSEC/DAY
71465 275/2330 04-283 OBC/CLK CLK ADJ FR -20.6 TO 0.0 MSEC/DAY
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 01 October 2004
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
-------- ----- ----------- ------
-None this report period.
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
-------- ----------- ------
-None this report period.
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 01 October 2004
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ----- -----------------
71374 269/2204 CLK ADJ FR -20.6 TO 0.0 MSEC/DAY
71449 274/2118 CLK ADJ FR 0.0 TO -20.6 MSEC/DAY
71465 275/2335 CLK ADJ FR -20.6 TO 0.0 MSEC/DAY
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 01 October 2004
24 September 2004 - GMT Day 268 Beta = 04.4 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 -700.0/-700.0 -3.36/-4.94 +00.0/+00.0 07.0 0.00 0.0
2 00.0/-50.4 +4.83/+3.87 +27.2/-12.0 25.9 1.02 19.3
3 11.2/-89.6 +3.87/+2.60 +26.8/-12.0 25.8 1.01 18.9
01 November 2004 - GMT Day 275 Beta = 28.6 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 -700.0/-700.0 -3.05/-3.68 +00.0/+00.0 07.0 0.00 0.0
2 00.0/-44.8 +5.79/+5.15 +27.2/-13.6 26.9 1.02 19.7
3 -16.8/-78.4 +4.51/+3.87 +26.0/-13.2 26.7 1.02 19.0
V/T LEVEL CHANGES
ORB# AOS TIME LVL FR-TO
----- -------- ---------
-None this report period.
APPENDIX E
UARS ANOMALY REPORTS
Week ending 01 October 2004
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
04-063 FOT 71416 272/1647 30832 Generic Late Acquisition
04-063 FOT 71433 273/2000 30832 Generic Late Acquisition
04-064 FOT 71435 273/2302 37111 Data Dropout / bad weather
04-065 FOT 71452 275/0256 37125 Bad Weather at GUAM
04-063 FOT 71463 275/1940 30832 Generic Late Acquisition
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
04-061 FOT 71286 264/0156 36997 Data Dropout
04-062 FOT 71288 264/0418 36998 Data Dropout
ANOMALY INVESTIGATION REPORTS GENERATED
AIR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
04-063 FOT See List Sept 04 30832 Generic Late Acquisition
04-064 FOT 71435 273/2302 37111 Data Dropout / bad weather
04-065 FOT 71452 275/0256 37125 Bad Weather at GUAM